# ID-Guard: A Universal Framework for Combating Facial Manipulation via Breaking Identification

### v1

@@ Abstract
The misuse of deep learning-based facial manipulation poses a potential threat to civil rights. To prevent this fraud at its source, proactive defense technology was proposed to disrupt the manipulation process by adding invisible adversarial perturbations into images, making the forged output unconvincing to the observer. However, their non-directional disruption of the output may result in the retention of identity information of the person in the image, leading to stigmatization of the individual. In this paper, we propose a novel universal framework for combating facial manipulation, called ID-Guard. Specifically, this framework requires only a single forward pass of an encoder-decoder network to generate a cross-model universal adversarial perturbation corresponding to a specific facial image. To ensure anonymity in manipulated facial images, a novel Identity Destruction Module (IDM) is introduced to destroy the identifiable information in forged faces targetedly. Additionally, we optimize the perturbations produced by considering the disruption towards different facial manipulations as a multi-task learning problem and design a dynamic weights strategy to improve cross-model performance. The proposed framework reports impressive results in defending against multiple widely used facial manipulations, effectively distorting the identifiable regions in the manipulated facial images. In addition, our experiments reveal the ID-Guard’s ability to enable disrupted images to avoid face inpaintings and open-source image recognition systems.
@@ I Introduction
The spread of false information in communities has been a longstanding concern, presenting a potential threat to civil rights and social security. In recent years, the advancement and deployment of generative deep neural networks (DNNs) have exacerbated this issue, with facial manipulation serving as a notable example. This technology enables end-to-end manipulation of facial attributes or identity of images/videos. Malicious actors, for instance, exploit forged images to generate and circulate misleading news [1, 2] or perpetrate online fraud [3]. Although the re-training of these methods is proved challenging due to high computing power requirements and technical thresholds, users can easily download accessible pre-trained models from open-source platforms such as Github  [^1] https://github.com, Hugging Face  [^2] https://huggingface.com, and TensorFlow Hub  [^3] https://www.tensorflow.org to implement the forgery [4]. This greatly reduces the cost of creating fake examples, thereby expediting the proliferation of false information on social media. Therefore, an urgent need exists to develop proactive and efficient defense mechanisms.
@@
Figure 1: Illustration of the impact of malicious propagation of facial manipulation samples. Fakes will lead to rumor spreading and insufficient distortion of faces by traditional defense methods will cause face stigmatization. Our method disrupts the observer’s identification of the identity in the sample and thus adequately protects the individual’s rights.
To mitigate the aforementioned threats, substantial research efforts have been directed toward proactive defense mechanisms against facial manipulation in recent times. Unlike passive detection methods [5, 6, 7, 8, 9], proactive defense algorithms [10, 11, 12, 4, 13, 14, 15, 16, 17, 18, 19] are crafted to thwart fraudulent activities at their source. These algorithms induce visual distortions in facial manipulation outputs by introducing imperceptible adversarial perturbations [20] into face images. However, the disruption caused by existing algorithms to manipulated images is often non-directional, leading to the preservation of individuals’ identifiable information in the disrupted images. When stubborn malicious users insist on uploading these incompletely disrupted forged images to social media, concerns such as face stigmatization may arise [16]. As illustrated in Fig. 1, applying the traditional proactive defense algorithm mentioned above to safeguard photographs of the renowned movie star Jackie Chan still allows ordinary Internet users to discern his identity readily. Insufficient distortion of facial images makes the disrupted output a “spoof”, leaving the victim’s reputation and rights inadequately protected. Notably, as shown in Fig. 2, above insufficiently disrupted facial examples also face two potential threats: 1) The residual identifiable information in the samples makes them still likely to be recognized by commercial facial recognition systems. The stigmatization problem is aggravated by the fact that some entertainment applications automatically identify and push images of celebrities; 2) Some malicious forgers with rich technical capabilities may restore the fake images that are not seriously distorted to continue to commit fraud.
@@
Figure 2: Illustration of potential threats to the insufficiently disrupted facial example. Challenges come primarily from commercial face recognition systems and face inpainting algorithms.
@@
To address the above concerns, in this paper, we propose a proactive defense framework called ID-guard. This framework requires only a single forward pass of an image reconstruction network to produce the universal adversarial perturbation to combat multiple open-source facial manipulation algorithms, as shown in Fig. 3. To destroy the identifiable semantic information in forged images and prevent spoofing of individuals, a novel Identity Destruction Module (IDM) is introduced. IDM enables the generated protective perturbations to focus on disrupting key regions of the individual’s face in an image to prevent an observer from recognizing his/her identity.
@@
In addition, the universality of the adversarial perturbations is crucial in practical application scenarios [12] due to the facial manipulation methods used by the forger are unknown and uncontrollable. To improve the cross-model universality of perturbations, a dynamic weighting strategy is proposed during the training of the perturbation generator. Specifically, the robustness of different facial manipulations against attacks varies due to differences in model structure and complexity. Simply treating the adversarial loss of attacking different models as equally weighted when training the perturbation generator will result in the produced perturbations being biased towards facial manipulations that are easy to attack. Therefore, we consider the attacks against different models as a multi-task learning problem, which enables the loss weights to be dynamically adjusted during the training process to achieve well-balanced cross-model performance. A gradient prior perturbation strategy is also introduced to guide the perturbation generator to improve the stability of training and accelerate convergence.
@@
As expected and will be verified by experiments, the proposed ID-guard can effectively distort recognizable regions in facial images forged by various open-source manipulation algorithms, thereby preventing observers and face recognition systems from recognizing their identities and circumventing face inpaintings. In summary, the contributions of our work are summarized as follows:
@@
    We propose a novel general adversarial perturbation generation framework to prevent facial manipulation from stigmatizing individuals, called ID-guard. Only a single forward pass of the generator is required to quickly produce perturbations that can disrupt various facial manipulations.
@@
    To fully disrupt the manipulation results so that the identities of individuals in the images cannot be identified, an Identity Destruction Module is proposed to guide the generated perturbations to focus on distorting identity-related semantic information. This module can further alleviate concerns about commercial face recognition systems and image inpainting algorithms.
@@
    To improve the universality of the generated adversarial perturbations, we implement an attack on multiple facial manipulations by solving a multi-task learning problem and designing a dynamic weighting strategy. To improve the stability of the generator, we introduce a gradient prior perturbation strategy.
The remainder of this paper is organized as follows: Related works on Deepfakes and proactive defense are reviewed in Section II. Section III presents the details of our method. The experimental results and analysis are provided in Section IV, followed by the conclusion in Section V.
Figure 3: The accessible pre-trained models can be easily downloaded from open-source platforms to implement forgeries. For a given image, the proposed ID-guard can generate universal perturbations for defense against multiple open-source facial manipulations through a single forward propagation of an image reconstruction network.
@@ II-A Facial Manipulation
Facial manipulation aims at constrained modification of the face of a given image/video to present the desired visual content, such as identity, expression, age, hair color, etc. of the character. In recent years, thanks to the great success of Generative Adversarial Networks (GANs) in image synthesis, many GAN-based algorithms [21, 22, 23, 24, 25] with different designs and constraints have been proposed to solve facial manipulation [26]. Some contributors choose to open source their work on the public platform, including pre-training models and running scripts, which greatly lowers the threshold for users to produce high-quality and high-fidelity fake images/videos.
II-B Facial Manipulation Disruption
Some recent works have achieved proactive defense against facial manipulation by injecting adversarial perturbations into the image. Ruiz et al.[10] and Yeh et al.[11] achieved disruption of the facial manipulation by deriving gradient-based adversarial perturbations on the target model. Works including [18] and [19] have significantly improved the robustness of adversarial perturbations for protecting personal images. However, the common drawbacks of these methods are higher computational overhead and lower work efficiency, due to they involve solving the model-specific perturbation for each image. Methods represented by [12, 4, 17] and [15] have studied cross-model universal perturbations, which to some extent alleviate overhead issues and practicality limitations. However, they ignore the fact that different facial manipulation models have differences in robustness and gradient direction, and thus the performance of the produced generic perturbations is uneven across models. Furthermore, as discussed above, these algorithms do not consider the problem of face stigmatization due to unconstrained disruptions. Zhai et al. [16] proposed a method to embed specific warning patterns on the generated fake images to solve this problem. Unlike them, the proposed ID-guard directly distorts the facial recognition area of fake images.
@@ II-C Multi-task Learning
One of the effective routes to achieve multi-task learning is to dynamically weight the losses of different tasks according to their learning stages or the difficulty of learning. Sener et al. [27] pointed out that multi-task learning can be regarded as a multi-objective optimization problem, aiming to find the Pareto optimal solution to optimize the performance of multiple tasks. A representative method that has been proven effective and widely used is the multiple gradient descent algorithm (MGDA) [28]. Some heuristic works [29, 30, 31, 32, 33] measured the difficulty of a task based on the order of magnitude or change rate of the loss value, and then dynamically adjusted the weights of different tasks to obtain balanced performance. In this work, we further explore the potential of integrating multi-task learning strategies into across-model universal perturbation generation.
@@ III Methodology
In this section, the specific design and implementation details of the proposed ID-guard framework are elaborated. For clarity, we first introduce the overview of the framework and the definition of notations.
III-A Preliminaries
III-A1 Adversarial Attacks against Facial Manipulations
Given a natural image $x\in\mathbb{R}^{3\times H\times W}$, the pre-trained facial manipulation model $\mathcal{M}$ translates it into a forgery $y=\mathcal{M}(x)$. The goal of the defender is to find a small adversarial perturbation $\delta$ that makes the facial manipulation model fail to manipulate the perturbed adversarial image $x_{adv}=x+\delta$. The solution to the $\delta$ can be formulated as a maximization optimization problem as follows
$\begin{split}&\max_{\eta}\mathcal{D}(\mathcal{M}(x),\mathcal{M}(x+\delta)),\\ &s.t.\quad\|\delta\|_{\infty}\leq\epsilon,\end{split}$ (1)
where $\mathcal{D}$ is the metric of the distance between images and its design is our focus and will be introduced in subsequent sections. $\epsilon$ is the infinite norm bound used to restrict the perturbation. The existing mainstream adversarial perturbation derivation patterns can be roughly divided into two categories: gradient-based methods [34, 35] and generation-based methods[36, 37]. Compared to the former, the latter is more flexible and faster, so we choose the generation-based approach to drive the perturbation derivation of the proposed ID-guard. Specifically, we train a perturbation generator $\mathcal{G}$ with the following optimization objective to produce the desired adversarial perturbation:
$\begin{split}&\max_{{\theta}_{\mathcal{G}}}\mathbb{E}(\mathcal{D}(\mathcal{M}(x),\mathcal{M}(x+\mathcal{G}(x)))),\\ &s.t.\quad\|\mathcal{G}(x)\|_{\infty}\leq\epsilon\end{split}$ (2)
$\mathcal{G}$ is designed as a Resnet architecture [38], and the advantages of this design will be discussed in subsequent experiments.
III-A2 Cross-Model Universal Perturbations
We hope that the adversarial perturbation generated for a certain image can be effective for a set of open-source pre-trained facial manipulation models $\mathcal{S}_{\mathcal{M}}=\left\{\mathcal{M}_{1},\mathcal{M}_{2},\dots,\mathcal{M}_{N}\right\}$. Hence, the optimization objective in Eq. (2) can be rewritten as:

### v2

@@ Abstract
The misuse of deep learning-based facial manipulation poses a significant threat to civil rights. To prevent this fraud at its source, proactive defense has been proposed to disrupt the manipulation process by adding invisible adversarial perturbations into images, making the forged output unconvincing to observers. However, the non-specific disruption against the output may lead to the retention of identifiable facial features, potentially resulting in the stigmatization of the individual. This paper proposes a universal framework for combating facial manipulation, termed ID-Guard. Specifically, this framework operates with a single forward pass of an encoder-decoder network to produce a cross-model transferable adversarial perturbation. A novel Identity Destruction Module (IDM) is introduced to degrade identifiable features in forged faces. We optimize the perturbation generation by framing the disruption of different facial manipulations as a multi-task learning problem, and a dynamic weight strategy is devised to enhance cross-model performance. Experimental results demonstrate that the proposed ID-Guard exhibits strong efficacy in defending against various facial manipulation models, effectively degrading identifiable regions in manipulated images. It also enables disrupted images to evade facial inpainting and image recognition systems. Additionally, ID-Guard can seamlessly function as a plug-and-play component, integrating with other tasks such as adversarial training.
@@ I Introduction
The spread of false information in communities has long been a major concern, posing a potential threat to civil rights and social security. The rapid advancement and widespread deployment of generative deep neural networks (DNNs) have further intensified this issue, with facial manipulation emerging as a prominent example. This technology enables end-to-end manipulation of facial attributes or identities of images and videos. Malicious actors, for instance, leverage forged images to fabricate and disseminate misleading news [7, 8] or engage in online fraud [6]. Although re-training these models remains challenging due to substantial computational demands and technical barriers, pre-trained models are readily available on open-source platforms such as GitHub  [^1] https://github.com, Hugging Face  [^2] https://huggingface.com, and TensorFlow Hub  [^3] https://www.tensorflow.org, enabling users to easily execute forgeries [9]. This significantly lowers the barrier to generating fake content, thereby accelerating the spread of misinformation on social media. Consequently, there is an urgent need to develop effective and proactive defense mechanisms.
@@
Figure 1: Illustration of the impact of malicious propagation of facial manipulation samples. Fakes will lead to rumors spreading, and insufficient distortion of faces by traditional defense methods will cause face stigmatization. Our method disrupts the observer’s identification of the identity in the sample and thus adequately protects the individual’s rights.
In response to these threats, significant research efforts have recently focused on developing proactive defense mechanisms against facial manipulation. Unlike passive detection methods [18, 19, 20, 56, 57], proactive defense algorithms [12, 23, 11, 9, 24, 13, 14, 15, 25, 27, 16] are designed to counteract fraudulent activities at their origin. However, these distortions suffer from several critical limitations: 1) they fail to completely obscure personally identifiable features, allowing identity-relevant information to persist, as the distortions produced by unconstrained adversarial perturbations appear in indeterminate locations rather than effectively covering the entire face; 2) they introduce random and unstructured artifacts, such as shadows, background distortions, and deformed facial regions, resulting in unnatural and visually unacceptable appearances. Consequently, if persistent malicious users insist on uploading these non-specifically disrupted images to social media, and these images are maliciously disseminated, concerns such as facial stigmatization may arise, posing ethical and reputational risks for the individuals depicted [15]. Especially for public figures, their identities remain discernible even after defense mechanisms are applied. For example, as illustrated in Fig. 1, when the above-mentioned traditional proactive defense is applied to protect a photo of the famous movie star Jackie Chan, his identity remains recognizable from the stigmatized forged output.
Furthermore, as illustrated in Fig. 2, inadequately disrupted facial images are also exposed to two potential threats: 1) The remaining identifiable information in these images increases the likelihood of their recognition by commercial facial recognition systems. This exacerbates the issue of stigmatization, as certain entertainment applications automatically detect and promote images of celebrities; 2) Technically adept malicious forgers may restore forged images that have not been significantly distorted by facial inpainting, enabling continued fraudulent activities.
@@
Figure 2: Illustration of potential threats to the insufficiently disrupted facial example. Challenges come primarily from commercial face recognition systems and facial inpainting algorithms.
Figure 3: The publicly available pre-trained models can be easily downloaded from open-source platforms to implement forgeries. For a given image, the proposed ID-Guard can generate transferable perturbations for defense against multiple open-source facial manipulations through a single forward propagation of an image reconstruction network.
@@
To address the above concerns, in this paper, we propose a proactive defense framework, ID-Guard. The framework generates transferable adversarial perturbations through a single forward pass of an image reconstruction network, effectively countering multiple open-source facial manipulation algorithms, as shown in Fig. 3. To eliminate identifiable semantic information in forged images and prevent identity spoofing, a novel Identity Destruction Module (IDM) is incorporated. The IDM disrupts identity-related features in a structured manner rather than introducing random or uncontrolled distortions, thereby ensuring that the individual’s identity remains unrecognizable.
@@
The transferability of adversarial perturbations is crucial in practical applications, as the facial manipulation methods employed by forgers are often unknown and uncontrollable. To address this, a dynamic weight strategy is introduced during the training of the perturbation generator. Specifically, the robustness of different facial manipulation models varies due to differences in model architecture and method design. Assigning equal weights to the adversarial loss of different models during perturbation generator training may cause the produced perturbations to be biased toward easily attacked facial manipulations, thereby degrading overall performance. Thus, adversarial attacks on different models are formulated as a multi-task learning problem, allowing the loss weights to be dynamically adjusted during training to achieve well-balanced cross-model performance. Additionally, a gradient prior perturbation strategy is introduced to enhance training stability and accelerate convergence.
@@
As anticipated and confirmed through experiments, the proposed ID-Guard effectively distorts identifiable regions in facial images manipulated by various open-source models, thereby preventing observers and face recognition systems from identifying individuals and bypassing facial inpainting techniques. Additionally, we demonstrate that ID-Guard can function as a plug-and-play module in adversarial training for facial manipulation models, thereby enhancing their adversarial robustness. In summary, the contributions of this work are summarized as follows:
@@
    We propose a novel general adversarial perturbation generation framework, termed ID-Guard, to prevent facial manipulation from stigmatizing individuals. A single forward pass of the generator suffices to produce perturbations capable of disrupting various facial manipulations. Moreover, this framework can be leveraged for seamless integration with other tasks.
@@
    To ensure complete disruption of manipulated images, preventing the identification of individuals, an Identity Destruction Module (IDM) is introduced. The IDM guides the generated perturbations to target identity-related semantic features, thereby mitigating concerns regarding commercial face recognition systems and image inpainting algorithms.
@@
    To improve the transferability of the generated adversarial perturbations, we implement attacks against multiple facial manipulations by solving a multi-task learning problem and designing a dynamic weight strategy. To improve the stability of the generator, a gradient prior perturbation strategy is introduced to improve the generator’s stability.
The remainder of this paper is organized as follows: Related works on facial manipulation and proactive defense are reviewed in Section II. Section III presents the details of our method. The experimental results and analysis are provided in Section IV, followed by the conclusion in Section V.
@@ II-A Facial Manipulation
Facial manipulation refers to the controlled modification of facial attributes in a given image or video to generate the desired visual content, including identity, expression, age, and hair color. In recent years, leveraging the remarkable success of Generative Adversarial Networks (GANs) in image synthesis, numerous GAN-based algorithms [1, 2, 3, 4, 5] featuring diverse architectures and constraints have been developed to facilitate facial manipulation [22]. Some researchers have opted to release their work as open-source on public platforms, providing pre-trained models and executable scripts, thereby significantly reducing the technical barriers for users to generate high-quality, high-fidelity fake images and videos.
II-B Proactive Defense against Facial Manipulation
From a defense objective standpoint, generalized proactive defense methods can be divided into two categories: proactive forensics and proactive disruption. Proactive forensics involves embedding imperceptible watermarks or traceable markers into multimedia content to facilitate the identification of manipulated samples. Using these embedded elements, defenders can verify content authenticity and trace the origins of facial manipulations. In contrast, proactive disruption seeks to degrade the quality of facial manipulation outputs through the injection of adversarial perturbations, thereby misleading the generative model. By distorting the generated results, proactive disruption effectively diminishes the realism and credibility of forged images.
II-B1 Proactive Forensics
Proactive forensic techniques aim to embed identifiable patterns into images to facilitate fake detection and manipulation provenance tracking. Early approaches, such as FaceGuard [58] and Faketagger[59], detected forged examples by embedding watermarks into real images and verifying their integrity upon retrieval, but lacked structured tracking mechanisms. To provide identity source tracking, Zhao et al. [63] embed watermarks as anti-Deepfake labels into facial identity features, enabling fake detection by verifying the presence of the label. These works mainly focus on authenticity detection, but cannot pinpoint the tampered region. To improve detection and localization, Asnani et al proposed a proactive embedding framework [60], later refining it into MALP [61], which integrates attention mechanisms to achieve fine-grained manipulation localization. PADL [62] further enhanced robustness by combining perturbation-based defenses with detection and localization strategies. Zhao et al. [64] embedded a semi-fragile watermark in the original image. Once counterfeited, the tampered regions can be located by comparing retrieved and original watermarks. The limitation of proactive forensics is that it preserves the integrity of forged samples, whereas proactive disruption directly degrades or nullifies the forgery. Therefore, in this paper, we focus on proactive disruption methods.
II-B2 Proactive Disruption
Recent studies have explored proactive disruption against facial manipulation by injecting adversarial perturbations into images. Ruiz et al.[12] and Yeh et al.[23] disrupted facial manipulation by deriving gradient-based adversarial perturbations on target models. Works including [27] and [16] have significantly improved the robustness of adversarial perturbations in protecting personal images. However, due to structural and design differences across various facial manipulation models, adversarial perturbations crafted for a specific model often exhibit poor transferability to other models, limiting their defense applicability. To address this, works such as [13] and [10] generated adversarial perturbations by attacking a surrogate model and transferred them to an inaccessible model. However, the significant structural differences between face manipulation models limit their effectiveness. A more widely adopted approach is model ensembling. Representative methods such as [11, 9, 25] and [14] have explored cross-model transferable adversarial perturbations based on this paradigm, which enhance defense effectiveness against various facial manipulation models to a certain extent. However, these approaches overlook the fact that different facial manipulation models exhibit variations in adversarial robustness and gradient optimization. As a result, the effectiveness of the generated transferable perturbations is inconsistent across models, leading to an overall decline in defense performance. This is one of the issues that this paper focuses on. Additionally, as noted earlier, these methods do not account for the issue of facial stigmatization caused by unconstrained perturbations. Zhai et al. [15] addressed this problem by embedding specific warning patterns into generated fake images. Unlike them, the proposed ID-Guard directly distorts the facial recognition area of fake images.
@@ II-C Multi-task Learning
One of the effective routes to achieve multi-task learning is to dynamically weight the losses of different tasks according to their learning stages or the difficulty of learning. Sener et al. [29] pointed out that multi-task learning can be regarded as a multi-objective optimization problem, aiming to find the Pareto optimal solution to optimize the performance of multiple tasks. A representative method that has been proven effective and widely used is the multiple gradient descent algorithm (MGDA) [30]. Some heuristic works [31, 32, 34, 33, 35] measured the difficulty of a task based on the order of magnitude or change rate of the loss value, and then dynamically adjusted the weights of different tasks to obtain balanced performance. In this work, we further explore the potential of integrating multi-task learning strategies into across-model transferable perturbation generation.
@@ III Methodology
In this section, the specific design and implementation details of the proposed ID-Guard framework are elaborated. For clarity, we first introduce an overview of the framework and a definition of notation.
III-A Overview
III-A1 Facial Manipulation
Facial manipulation can be regarded as an image translation task, aiming to transform a given original example into a target manipulated example. Specifically, given an original face image $x\in\mathbb{R}^{3\times H\times W}$, the facial manipulation model $\mathcal{M}$ leverages the specified target attribute or identity $a$ to map it to a forged image $y$. The process is formulated as:
$\begin{split}y=\mathcal{M}(x,a).\end{split}$ (1)
Since end-to-end facial manipulation models employ different methods for embedding attributes or identity information, we simplify the manipulation process as $y=\mathcal{M}(x)$. where the face in the manipulated image $y$ retains the identity of the person in the original image $x$ while exhibiting the specified features of attribute $a$.
III-A2 Adversarial Perturbation against Manipulation
To proactively disrupt facial manipulation, the defender’s objective is to generate an imperceptible adversarial perturbation $\delta$ such that the target facial manipulation model fails to map the adversarial image $x_{adv}=x+\delta$ to an acceptable manipulated output. The optimization of $\delta$ can be formulated as a maximization problem:
$\begin{split}&\max_{\delta}\mathcal{D}(\mathcal{M}(x),\mathcal{M}(x+\delta)),\\ &s.t.\quad\|\delta\|_{\infty}\leq\epsilon,\end{split}$ (2)
where $\epsilon$ is the infinite norm bound used to restrict the perturbation, and $\mathcal{D}$ is the distance metric between the original forged image $\mathcal{M}(x)$ and disrupted forged image $\mathcal{M}(x+\delta)$. Some existing works [12, 23] employ gradient-based adversarial attack algorithms[36, 37] to generate perturbations. However, these methods require multiple iterations for each perturbation generation, leading to high computational overhead. A more efficient approach is to train a perturbation generator $\mathcal{G}$, which can produce adversarial perturbations with a single forward pass during inference, i.e., $\delta=\mathcal{G}(x)$. The optimization problem for training $\mathcal{G}$ is defined as follows:
$\begin{split}&\max_{{\theta}_{\mathcal{G}}}\mathbb{E}(\mathcal{D}(\mathcal{M}(x),\mathcal{M}(x+\mathcal{G}(x)))),\\ &s.t.\quad\|\mathcal{G}(x)\|_{\infty}\leq\epsilon,\end{split}$ (3)
where ${\theta}_{\mathcal{G}}$ is the parameter of the perturbation generator. On one hand, previous state-of-the-art methods [12, 23, 10, 11] typically apply the Mean Squared Error (MSE) loss as a proxy for the distance metric $\mathcal{D}$. However, this approach often results in unstructured distortions in the manipulated image while preserving identifiable facial features, leading to facial stigmatization. On the other hand, the perturbations generated by these methods generally lack transferability and can only disrupt a single target model $\mathcal{M}$.
III-A3 ID-Guard Framework
To address the aforementioned issues, in this paper, we propose the ID-Guard framework, which aims to train a generator $\mathcal{G}$ with Resnet[51] architecture capable of producing the cross-model transferable adversarial perturbation, as shown in Fig. 4. This generator produces customized perturbations for given images, effectively defending against a set of pre-trained facial manipulation models $\mathcal{S}_{\mathcal{M}}=\left\{\mathcal{M}_{1},\mathcal{M}_{2},\dots,\mathcal{M}_{N}\right\}$. Following the model ensemble paradigm, for all accessible models $\mathcal{M}_{k}$, Eq.(3) in our proposed framework can be rewritten as: